# Higgs boson production at $\mu^{+}\mu^{+}$ colliders

### v1

@@ 1 Introduction
Recently, muon colliders based on ultra-slow muon technology Nagamine:1995zz have been proposed Hamada:2022mua . This is because ultra-slow muons from laser-ionized muonium, a $\mu^{+}e^{-}$ bound state, can be used to create $\mu^{+}$ beams with excellent emittances, i.e., very collimated beams that facilitate high luminosities. As such, the corresponding proposal is to build on this technology, which was developed for the muon $g-2$/EDM experiment at J-PARC Abe:2019thb , leading to an estimated luminosity seemingly good enough for Higgs factories based on $\mu^{+}e^{-}$ and $\mu^{+}\mu^{+}$ colliders.
@@
At such colliders, even though there are no $s$-channel annihilation processes that are present in $\mu^{+}\mu^{-}$ or $e^{+}e^{-}$ colliders, Higgs boson production can still occur via vector boson fusion processes, which dominate over $s$-channel annihilations at energies beyond $\mathcal{O}(1)$ TeV. This is due to a logarithmic enhancement of the cross section for vector boson fusion as a function of center-of-mass energy. At $\mu^{+}e^{-}$ colliders, while $W$ boson fusion is possible, the center-of-mass energy is limited by the beam energy of the electron. On the other hand, for $\mu^{+}\mu^{+}$ colliders, one can consider, e.g., $\mathcal{O}(10)$ TeV beam energies, but there is no $W$ boson fusion process at leading order since both of the antimuons can only emit positively charged $W$ bosons. Although $Z$ boson fusion is possible at leading order since it is independent of the muon charge, its cross section is about an order of magnitude smaller than $W$ boson fusion at $\mu^{+}\mu^{-}$ colliders due to an unfortunate suppression of the coupling between the $Z$ boson and leptons. This motivates us to look beyond the leading order.
@@
In this paper, we thus investigate $W$ boson fusion at $\mu^{+}\mu^{+}$ colliders at higher order in perturbation theory. The pertinent process, which we show in Fig. 1, begins with emission of a photon or $Z$ boson from one of the antimuons, followed by the photon or $Z$ boson splitting into a $W^{+}W^{-}$ pair. The $W^{-}$ from this splitting then collides with the $W^{+}$ emitted from the other antimuon to produce a Higgs boson. For large center-of-mass energies $\sqrt{s}$, we find that this sub-leading-order process involves a $(\log s)^{3}$ factor, in contrast to the single $\log s$ appearing in the leading-order $W$ boson fusion at $\mu^{+}\mu^{-}$ and $e^{+}e^{-}$ colliders AlAli:2021let ; ILC:2013jhg . Due to this large enhancement, the cross section for Higgs boson production at $\mu^{+}\mu^{+}$ colliders becomes comparable to that at $\mu^{+}\mu^{-}$ colliders for $\mathcal{O}(10)$ TeV beam energies.
@@ 2.1 Triple logarithms in the Higgs boson production
we observe a rapid growth of the cross section as a function of the collider energy. Furthermore, the cross section of the $Z$ boson fusion process has an unfortunate suppression factor in the numerator. Because of this, at $\sqrt{s}\sim\mathcal{O}(10)$ TeV, the sub-leading, $\gamma$-mediated $W$ boson fusion cross section becomes much larger than for the leading-order $Z$ boson fusion process. Even when compared to the cross section of the leading-order $W$ boson fusion process at $\mu^{+}\mu^{-}$ or $e^{+}e^{-}$ colliders Cahn:1983ip ; Cahn:1984tx ,
@@ 2.3 Numerical calculation with MadGraph
We furthermore overlaid the cross section of the leading-order $W$ boson fusion process at $\mu^{+}\mu^{-}$ colliders, where we do not assume beam polarizations. We see that at energies of $10$ TeV or above, the cross section of the higher-order process with polarized beams becomes about half as large as the leading one at $\mu^{+}\mu^{-}$ colliders. See also Tab. 1, where we list the cross sections for representative collider energies. Therefore, the disadvantage due to missing $W$ boson fusion at the leading order is remedied at the next-to-leading order at $\sim 10$ TeV energies. Importantly, this large contribution to Higgs production is also present for $\mu^{+}\mu^{-}$ and $e^{+}e^{-}$ colliders. This will thus add a new significant process at any such lepton collider.
Lastly, we also show in Fig. 2 the cross section of $Z$ boson fusion at $\mu^{+}\mu^{+}$ colliders for $P_{\mu^{+}}=+0.8$ and $P_{\mu^{+}}=0$. We see that the cross section of the sub-leading $W$ boson fusion process becomes larger than for the leading $Z$ boson fusion beyond a few TeV.
The observed large correction to $W$ boson fusion at the next-to-leading order may indicate that we are about to lose perturbativity. We thus investigate higher-order processes with additional powers of $\log s$, which are, e.g., $\mu^{+}\mu^{+}\to\mu^{+}\mu^{+}W^{+}W^{-}h$, $\mu^{+}\mu^{+}\to\mu^{+}\overline{\nu}_{\mu}W^{+}\,\gamma/Z\,h$ and $\mu^{+}\mu^{+}\to\overline{\nu}_{\mu}\overline{\nu}_{\mu}W^{+}W^{+}h$. In particular, the first process has a parametric $(\log s)^{5}$ dependence, while the others have $(\log s)^{4}$, and we thus expect it to have the largest contribution at high energies among them. Using the method we describe in Sec. 3, we calculate the cross section of this process to be $\sigma\simeq 14$ fb at $\sqrt{s}=20$ TeV for $P_{\mu^{+}}=+0.8$, which is notably smaller than the cross section $\sigma\simeq 600$ fb of the lower-order process $\mu^{+}\mu^{+}\to\mu^{+}\overline{\nu}_{\mu}W^{+}h$. Therefore, at least for Higgs boson production at $\mathcal{O}(10)$ TeV energies, the higher-order processes are not yet in the non-perturbative regime. Nevertheless, in the analysis of actual events, one should include the higher-order processes since they may contribute to the signals defined by lower-order ones.
@@ 3.1 Low-$p_{\rm T}$ cross sections from EPA
$f^{p_{\rm T}^{\rm(cut)}}_{\gamma/\mu}\!\!\!(x)$ $=$ $\frac{\alpha}{2\pi}\left[-\frac{2(1-x)}{x}\frac{\left(p_{\rm T}^{\rm(cut)}\right)^{2}}{\left(p_{\rm T}^{\rm(cut)}\right)^{2}+m_{\ell}^{2}x^{2}}+\frac{1+(1-x)^{2}}{x}\log\frac{\left(p_{\rm T}^{\rm(cut)}\right)^{2}+m_{\ell}^{2}x^{2}}{m_{\ell}^{2}x^{2}}\right]$ (21)

### v2

@@ 1 Introduction
Recently, muon colliders based on ultra-slow muon technology Nagamine:1995zz have been proposed Hamada:2022mua . This is because ultra-slow muons from laser-ionized muonium, a $\mu^{+}e^{-}$ bound state, can be used to create $\mu^{+}$ beams with excellent emittances, i.e., very collimated beams that facilitate high luminosities. As such, the corresponding proposal is to build on this technology, which was developed for the muon $g-2$/EDM experiment at J-PARC Abe:2019thb . Thus, under various assumptions for the parameters of the proposed collider facility, the luminosity is estimated to be $\sim 5.7\times 10^{32}$ cm-2s-1 for a $\mu^{+}\mu^{+}$ collider with center-of-mass energy 2 TeV, and $\sim 4.6\times 10^{33}$ cm-2s-1 for a $\mu^{+}e^{-}$ collider with center-of-mass energy 346 GeV Hamada:2022mua . For higher energies, the improvement in the emittance as well as the larger boost factor further enhance the luminosity. In any case, this estimate suggests that these colliders are realistic options for Higgs boson factories.
@@
At such $\mu^{+}$-based colliders, even though there are no $s$-channel annihilation processes that are present in $\mu^{+}\mu^{-}$ or $e^{+}e^{-}$ colliders, Higgs boson production can still occur via vector boson fusion processes, which dominate over $s$-channel annihilations at energies beyond $\mathcal{O}(1)$ TeV. This is due to a logarithmic enhancement of the cross section for vector boson fusion as a function of center-of-mass energy. At $\mu^{+}e^{-}$ colliders, while $W$ boson fusion is possible, the center-of-mass energy is limited by the beam energy of the electron. On the other hand, for $\mu^{+}\mu^{+}$ colliders, one can consider, e.g., $\mathcal{O}(10)$ TeV beam energies, but there is no $W$ boson fusion process at leading order since both of the antimuons can only emit positively charged $W$ bosons. Although $Z$ boson fusion is possible at leading order since it is independent of the muon charge, its cross section is about an order of magnitude smaller than $W$ boson fusion at $\mu^{+}\mu^{-}$ colliders due to an unfortunate suppression of the coupling between the $Z$ boson and leptons. This motivates us to look beyond the leading order.
@@
In this paper, we thus investigate $W$ boson fusion at $\mu^{+}\mu^{+}$ colliders at higher order in perturbation theory111 Here, we mean “higher order” with respect to the leading-order $W$ boson fusion process at $\mu^{+}\mu^{-}$ colliders, i.e., $\mu^{+}\mu^{-}\to\overline{\nu}_{\mu}\nu_{\mu}h$, rather than higher-order corrections to some process at $\mu^{+}\mu^{+}$ colliders. . The pertinent process, which we show in Fig. 1, begins with emission of a photon or $Z$ boson from one of the antimuons, followed by the photon or $Z$ boson splitting into a $W^{+}W^{-}$ pair. The $W^{-}$ from this splitting then collides with the $W^{+}$ emitted from the other antimuon to produce a Higgs boson. For large center-of-mass energies $\sqrt{s}$, we find that this higher-order process involves a $(\log s)^{3}$ factor, in contrast to the single $\log s$ appearing in the leading-order $W$ boson fusion at $\mu^{+}\mu^{-}$ and $e^{+}e^{-}$ colliders AlAli:2021let ; ILC:2013jhg . Due to this large enhancement, the cross section for Higgs boson production at $\mu^{+}\mu^{+}$ colliders becomes comparable to that at $\mu^{+}\mu^{-}$ colliders for $\mathcal{O}(10)$ TeV beam energies.
@@ 2.1 Triple logarithms in the Higgs boson production
In Eq. (5), we observe three powers of logarithms increasing with the center-of-mass energy. These originate from collinear and soft divergences arising in the limit of massless muons, $W$ bosons, and Higgs bosons. In particular, the photon emission from one $\mu^{+}$ has an associated collinear divergence $\sim\log s/m_{\mu}^{2}$ and soft divergence $\sim\log s/s_{\rm min}$, stemming respectively from the logarithm and $1/x$ factor in the parton distribution function of Eq. (1). Similarly, the $W^{+}$ emitted by the other $\mu^{+}$ has an associated soft divergence $\sim\log s/s_{\rm min}$, but no collinear divergence since it is longitudinally polarized, and thus its parton distribution function in Eq. (3) only has a $1/x$ factor and no collinear logarithm. The second term $\sim(\log s/s_{\rm min})^{3}$ in the brackets of Eq. (5) arises because our chosen factorization scale depends on momentum fractions, and would for instance not appear for $\mu_{f}^{2}=s$. Our choice of factorization scale, and hence the additional term, serves to improve the (numerical) accuracy of the parton distribution function approximation.
@@
we observe a rapid growth of the cross section as a function of the collider energy. Furthermore, the cross section of the $Z$ boson fusion process has an unfortunate suppression factor in the numerator. Because of this, at $\sqrt{s}\sim\mathcal{O}(10)$ TeV, the higher-order, $\gamma$-mediated $W$ boson fusion cross section becomes much larger than for the leading-order $Z$ boson fusion process. Even when compared to the cross section of the leading-order $W$ boson fusion process at $\mu^{+}\mu^{-}$ or $e^{+}e^{-}$ colliders Cahn:1983ip ; Cahn:1984tx ,
@@ 2.3 Numerical calculation with MadGraph
We furthermore overlaid the cross section of the leading-order $W$ boson fusion process at $\mu^{+}\mu^{-}$ colliders, where we do not assume beam polarizations. We see that at energies of $10$ TeV or above, the cross section of the higher-order process with polarized beams becomes about half as large as the leading one at $\mu^{+}\mu^{-}$ colliders. Therefore, the disadvantage due to missing $W$ boson fusion at the leading order is remedied by the higher-order process at $\sim 10$ TeV energies. Importantly, this large contribution to Higgs production is also present for $\mu^{+}\mu^{-}$ and $e^{+}e^{-}$ colliders. This will thus add a new significant process at any such lepton colliders.
Lastly, we show in Fig. 2 the cross section of $Z$ boson fusion at $\mu^{+}\mu^{+}$ colliders for $P_{\mu^{+}}=+0.8$ and $P_{\mu^{+}}=0$. We see that the cross section of the higher-order $W$ boson fusion process becomes larger than that of the leading $Z$ boson fusion beyond a few TeV.
For comparison, we list the cross sections for representative collider energies in Tab. 1. We computed these values using the same methods as the respective curves in Fig. 2, i.e., using MadGraph for $\sigma(\mu^{+}\mu^{-}\to\overline{\nu}\nu h)$ and $\sigma(\mu^{+}\mu^{+}\to\mu^{+}\mu^{+}h)$, and the full calculation with controlled theoretical uncertainties described in Sec. 3 for $\sigma(\mu^{+}\mu^{+}\to\mu^{+}\overline{\nu}W^{+}h)$. Note that in Tab. 1, we omit the theoretical uncertainties of the cross sections, which are of $\mathcal{O}(1)$ fb or smaller.
The observed enhancement due to the large logarithms in $W$ boson fusion at a higher order may indicate a breakdown of the perturbative expansion. If this is the case, one needs to resum higher-order corrections stemming from multiple soft and collinear emissions. Such resummation effects are encoded in electroweak parton distribution functions by solving a set of coupled differential equations, called the electroweak DGLAP equations Ciafaloni:2001mu ; Ciafaloni:2005fm ; Chen:2016wkt ; Bauer:2017isx ; Bauer:2017bnh ; Manohar:2018kfx ; Fornal:2018znf ; Han:2020uid ; Han:2021kes ; Azatov:2022itm ; Garosi:2023bvq ; Marzocca:2024fqb .
Here, to verify the validity of our fixed-order calculation, we evaluate the cross section of the process $\mu^{+}\mu^{+}\to\mu^{+}\mu^{+}W^{+}W^{-}h$. Parametrically, this process has a logarithmic enhancement factor of $\alpha^{5}\,(\log s)^{5}$, compared to $\alpha^{4}\,(\log s)^{3}$ for the process $\mu^{+}\mu^{+}\to\mu^{+}\overline{\nu}W^{+}h$. Thus, the expansion parameter is $\alpha\,(\log s)^{2}$, which is the largest correction factor we expect. The size of the $\mu^{+}\mu^{+}\to\mu^{+}\mu^{+}W^{+}W^{-}h$ cross section therefore serves as an indicator of the reliability of the perturbative expansion. Using the method we describe in Sec. 3, we obtain the cross section $\sigma\simeq 14$ fb at $\sqrt{s}=20$ TeV for $P_{\mu^{+}}=+0.8$, which is notably smaller than the cross section $\sigma\simeq 600$ fb of the lower-order process $\mu^{+}\mu^{+}\to\mu^{+}\overline{\nu}W^{+}h$. Therefore, at least for Higgs boson production at $\mathcal{O}(10)$ TeV energies, such higher-order processes are not yet in the non-perturbative regime. Note that we expect to lose perturbativity at $\mathcal{O}(100)$ TeV energies, where the logarithmic enhancement factor combined with the coupling constant, $\sim(\alpha/\pi)\log(s/m_{\mu}^{2})\log(s/s_{\rm min})$, becomes $\mathcal{O}(1)$. Nevertheless, in the analysis of actual events, one should include the higher-order processes since they may contribute to the signals defined by lower-order ones.
@@ 3.1 Low-$p_{\rm T}$ cross sections from EPA
$f^{p_{\rm T}^{\rm(cut)}}_{\gamma/\ell}\!\!\!(x)$ $=$ $\frac{\alpha}{2\pi}\left[-\frac{2(1-x)}{x}\frac{\left(p_{\rm T}^{\rm(cut)}\right)^{2}}{\left(p_{\rm T}^{\rm(cut)}\right)^{2}+m_{\ell}^{2}x^{2}}+\frac{1+(1-x)^{2}}{x}\log\frac{\left(p_{\rm T}^{\rm(cut)}\right)^{2}+m_{\ell}^{2}x^{2}}{m_{\ell}^{2}x^{2}}\right]$ (21)